OpenStudy (anonymous):

write the following ellipse in standard form: 2(x+4)²+3(y-1)²=24

OpenStudy (mathstudent55):

Divide both sides by 24

OpenStudy (anonymous):

@mathstudent55 so 2(x+4)²/24+3(y-1)²/24? in that case what would the answer be

OpenStudy (mathstudent55):

Now simplify each fraction on the left side. In the first one you have 2 in the numerator and 24 in the denominator. In the second one you have 3 in numerator and 24 in denominator. On right side you have 24/24 which is 1.

OpenStudy (mathstudent55):

2(x + 4)²/24 + 3(y - 1)²/24 = 24/24 (x + 4)²/12 + (y - 1)²/8 = 1

OpenStudy (anonymous):

thanks so so much!!!! i understand it so much better now

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

@mathstudent55 in this problem do you know how to find the sum of the focal radii?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The denominators are a^2 and b^2, respectively in a standard eq for an ellipse. The sum of the radial foci is a + b, or sqrt(12) + sqrt(8). But this needs to be simplified.